OpenStudy (anonymous):

4x+y=13 x+-y=2 ?x= ?

OpenStudy (anonymous):

I want to help you lead through this problem. Let's start by looking at the two equations, add the left and right sides and tell me what you get.

OpenStudy (anonymous):

17?

OpenStudy (anonymous):

Ok... got that...

OpenStudy (anonymous):

Let me try to illustrate this 4x+y=13 x -y= 2 See how I lined the problem up? It kind of looks like a giant equation. --------- 4x+x+y-y=15. We can add the left side and the right side. See what I did? Now let's combine the terms. 5x+0=15. The y's cancel out because y-y=0, we are left with... 5x=15 x=3. Now let's plug x=3 in to one of our equations x-y=2 3-y=2 3=y+2 3-2=y y=1.

OpenStudy (anonymous):

Thank you, that was very helpful!